In an attempt to get the last parking spot at West step by step pls
Answer #1

The car travelling at speed v sounds its horn. The sound waves approaching the wall are doppler shifted to a higher frequency f1 by the doppler formula:

f1 = (1 + v/c) f0 where c = speed of sound in air and f0 = frequency of horn at rest.

The waves are reflected but now the car is moving toward what appears to be the source of a sound of frequency f1 so there is again a upshift in frequency to f2

f2 = (1 +v/c) f1

f2 = (1 + v/c)^2 f0

Now since the driver hear 8 beats per second, it means the driver hears a frequency of 160 + 8 = 168 Hz

=> 168 = (1+v/c)^2*160

=> 168/160 = (1+v/c)^2

1.05 = (1+v/c)^2

1 + v/c = 1.0246

v/c = 0.0246

v = 0.0246*343 = 8.43 m/s